OpenStudy (anonymous):

@madiiharrell @taylorfun123@naguhty-girl

OpenStudy (anonymous):

what the question

OpenStudy (anonymous):

ok im looking at this problem one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the first i think

OpenStudy (anonymous):

r u sure i really dont want to fail and it says ally that applys

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so ur not sure

OpenStudy (anonymous):

do anyone know this problem?

OpenStudy (anonymous):

one sec im figuring it out just hold your horses

OpenStudy (anonymous):

ive done this problem befoire i just forgot the answers so hold on ill tell you when im done ok

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

what is one written as a fraction with 8 in the denominator. if you can do that you will have your answer

OpenStudy (anonymous):

but i dont knwo how to do it can u help me ?

OpenStudy (anonymous):

\[ 1 = \frac{ x }{ 8 }\]

OpenStudy (anonymous):

do taht equation right

OpenStudy (anonymous):

thanks Austin

OpenStudy (anonymous):

the answer is \[\frac{ 8 }{ 8 } = 1\]

OpenStudy (anonymous):

so is it teh first one

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then....

OpenStudy (anonymous):

can u help me understand it

OpenStudy (anonymous):

and 9+3 =12 which is 12/8

OpenStudy (anonymous):

which is the last one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its not teh last one then whta is it

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but let Austin help you i have another problem i have to figure out sorry

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

SO is it C

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its then one you have not said yet

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yes -_-

OpenStudy (anonymous):

ITS only b

OpenStudy (anonymous):

no D would work i guess

OpenStudy (anonymous):

r u sure u said no for d

OpenStudy (anonymous):

yeah i guess i got ahead of myself

OpenStudy (anonymous):

u cant help no more ?